# The Science of Data Collection: Insights from Surveys can Improve Machine Learning Models

### v1

@@ Abstract
Whether future AI models make the world safer or less safe for humans rests in part on our ability to efficiently collect accurate data from people about what they want the models to do. However, collecting high quality data is difficult, and most AI/ML researchers are not trained in data collection methods. The growing emphasis on data-centric AI highlights the potential of data to enhance model performance. It also reveals an opportunity to gain insights from survey methodology, the science of collecting high-quality survey data.
In this position paper, we summarize lessons from the survey methodology literature and discuss how they can improve the quality of training and feedback data, which in turn improve model performance. Based on the cognitive response process model, we formulate specific hypotheses about the aspects of label collection that may impact training data quality. We also suggest collaborative research ideas into how possible biases in data collection can be mitigated, making models more accurate and human-centric.
@@ 1 Introduction
AI and ML practitioners are not the only researchers who depend on data collected from humans to achieve their aims. Social scientists have long relied on survey data collected from human subjects to quantify the population, understand public opinion, and test hypotheses about human behavior. The methods used to collect survey data have been extensively studied and refined by researchers in the field of survey methodology, which draws on social and cognitive psychology to develop theories about how humans understand, process and respond to questions in surveys (Groves et al., 2009).
Because human-labeled data is central to all stages of the AI pipeline (Plank, 2022; Mazumder et al., 2023), it is surprising that so little research in the AI literature has used social science, and survey methodology in particular, to understand the actions and motivations of the humans behind the data generating process. Insights from social science will be critical to the development of more trustworthy and human-centric models: “if we want to train AI to do what humans want, we need to study humans” (Irving & Askell, 2019).
Collecting high quality data is difficult, as decades of research in survey methodology and recent high-profile failures in opinion polling (Sturgis et al., 2016; Kennedy et al., 2017; Clinton et al., 2021) demonstrate. We worry that many researchers collecting data to train, fine-tune, or reinforce AI and ML models are not trained in data collection. A recent paper lamented: “everyone wants to do the model work, not the data work” (Sambasivan et al., 2021). This paper introduces AI researchers to the community of scientists who want to do the data work and their insights into how to collect high quality data.
We first make the case that label collection is similar to survey data collection. Then, we explore what we learn when we think of labelers as survey respondents (Section 3) and as interviewers (Section 4). In Section 5, we join the call for greater transparency in label collection methods, offering lessons from surveys and statistical methods. Throughout the paper, we use the terms labels (and labelers) to refer generally to ML annotations, such as image bounding boxes, natural language understanding labels, model evaluation data, human feedback for reinforcement learning, and other types of training data (and data collectors).
@@ 2 How Labeling is Like a Survey
Surveys can be self-administered, like a web or a paper-and-pencil survey, or interviewer administered, like a telephone or face-to-face survey. A web survey presents a series of questions with, most often, closed answer choices. Superficially, the process of labeling observations for ML models often looks like a web survey. Labelers see one or more prompts and associated answer choices (Figure 1). As in surveys, the task can be factual (is there a bicycle in this photo?) or opinion based (which of these responses to the prompt is most helpful?). From this perspective, the labelers are survey respondents.
However, from a different perspective, labelers are interviewers. In a survey, respondents answer each question only once, but interviewers ask the same question dozens of times. Subtle differences in the way interviews ask questions introduce interviewer effects (Groves, 1989; Schnell & Kreuter, 2005). Like interviewers, labelers respond to a similar prompt many times, and subtle differences in how they label lead to labeler effects (Geva et al., 2019). In addition, labelers, like interviewers, are provided with training and feedback (Pustejovsky & Stubbs, 2012; Fort, 2016) and are usually compensated for their labeling work.
@@
Of course, label collection differs from survey data collection in important ways. Surveys usually ask people about themselves: their opinions, behaviors, characteristics. Labeling tasks more often involve passing judgment on an object outside of oneself (images, product reviews). However, the human feedback used in reinforcement learning may be more like a survey question: “which of the following responses is most relevant to the prompt?” Surveys ask many different questions and end when the questions run out; labeling tasks often consist of similar repeated observations, and labelers can choose when to stop.
@@
The goal of surveys and labeling tasks is also different. Surveys ask questions of a sample of selected persons to make inference about the population. For example, say we wished to know what proportion of the US adult population does not have health insurance. We select a sample of US adults and ask them whether they have health insurance. For each respondent, we have a yes or no response to the question. [^1] We are ignoring “I don’t know” and other types of responses. The proportion of respondents without insurance in the sample is an estimate of the proportion of the population without insurance. A survey often asks many questions on related topics. Our example survey might also ask “when was the last time you visited a health care professional?” and “have you received a flu vaccine this year?” Results are often reported at the question level although relationships between questions are also of interest: for example, are people with health insurance more likely to get a flu vaccine? The quality of the analysis rests on collecting correct responses from a representative sample of US adults.
@@
The goal in labeling work is not to estimate the percentage of the population that finds a given post offensive or detects a vehicle in a given image. [^2] Many studies collect only one or a few labels of each instance, or collapse multiple labels per instance, and thus cannot capture the perspective and diversity of a population, despite research showing that this diversity improves models (Basile et al., 2021; Sap et al., 2022) Instead, the goal in training is to learn patterns from the labeled data to predict labels for unseen observations. Thus, while it is not as important that labels reflect the population’s views at an instance level, across instances, the views of the population should be represented Mitchell et al. (2019).
3 Labelers as Respondents
We first consider what we can learn from the survey methods literature when we consider labelers as respondents.
3.1 Relevant Social Science Theories
Social scientists have developed several theories about how respondents understand survey questions and formulate answers to them.
@@ 3.1 Relevant Social Science Theories
Responding to a survey question involves several cognitive steps (Tourangeau et al., 2000; Tourangeau, 2018). The simplest version of the cognitive response process model has four steps:
@@
Ideally, a respondent proceeds through each step in order. However, at each step, the respondent can choose to repeat an earlier step. For example, considering the response options in the Mapping step may change the interpretation of the question (Comprehension) or trigger additional relevant information to come to mind (Retrieval)).
The cognitive response process model exposes why respondents sometimes give incorrect answers. At the Comprehension step, they may fail to understand the question or some of the words it uses. They may have a different understanding of some of the words than those who wrote the question. At the second step, respondents may fail to retrieve all relevant information. Some information may have been forgotten. At the third step, respondents may fail to put in the mental effort to bring together their understanding of the question with the retrieved information. At the fourth step, respondents may not find an answer choice that reflects their answer or they may edit the true answer to avoid revealing sensitive information.
@@
The full survey response process outlined above places a cognitive burden on respondents. Many resort to taking shortcuts, an approach called satisficing (Krosnick, 1991; Krosnick et al., 1996). For example, respondents may retrieve only the most recent relevant information from memory (recency bias) or choose the first reasonably correct answer choice.
Satisficing relates to the cognitive miser theory in psychology, which holds that people try to minimize cognitive effort (Fiske & Taylor, 1991; Kahneman, 2011). As predicted by satisficing theory, eye-tracking studies show that respondents do not read all options in multi-choice questions (Galesic et al., 2008), and shortcuts are more common as survey length increases (Galesic & Bosnjak, 2009). Respondents in survey tasks tend not to read provided instructions Brosnan et al. (2019) or click on provided definitions (Peytchev et al., 2010), especially when they think they know the concept that is asked about Roger et al. (2006).
Acquiescence, the tendency to say “yes” to yes/no questions, regardless of content, is another shortcut response technique (Knowles & Condon, 1999). Some respondents even deliberately give incorrect answers to later questions to reduce the burden of a survey, a phenomenon called motivated misreporting (Kreuter et al., 2011; Tourangeau et al., 2012; Eckman et al., 2014).
@@
Context effects concern how human perception is influenced by information perceived previously (Tversky & Kahneman, 1974; Strack, 1992). For example, a very tall person can make others seem shorter: a contrast effect. An unethical politician can make other politicians seem less ethical: an assimilation effect (Bless & Schwarz, 2010).
In surveys, attitudes and opinions are especially vulnerable to context effects, because respondents do not always have well-formed, fixed opinions that they retrieve from memory. Instead, they form opinions when asked for them, and this process can be shaped by context clues in the sentence, the response options, the instrument, or the previous questions (Zaller & Feldman, 1992; Schwarz, 2007).
@@ 3.2 Hypotheses about Label Quality
These theories about how respondents answer questions lead us to several hypotheses about the properties of the labeling task that may impact training data quality. The ML literature has investigated some of these hypotheses, but despite the isolated efforts mentioned below, there exist fundamental research gaps.
Observation Order
Theories about context effects suggest that the order in which instances or observations are presented influences the labels assigned. If a contrast effect is present, a very hateful social media post would make later posts seem less hateful than they otherwise would (for preliminary evidence of this phenomenon, see Beck et al., 2022). An order effect could also arise if labelers change their behavior over time. As they gain experience, they might become more accurate and conscientious, as suggested by Lee et al. (2022). Alternatively, they might become bored and engage in more satisficing over time, as suggested by the survey literature (Galesic & Bosnjak, 2009). Many research questions thus remain open, for instance, how and to what extent observation order impacts training data, what type of tasks are most impacted by order effects, and the implications of order effects for active learning and similar label collection approaches.
@@
Often ML researchers want to collect multiple labels about an observation from one labeler: for example, whether an image contains a cat, a dog, a person, or a vehicle. We can ask labelers to provide all labels at once, as in the left panel of Figure LABEL:fig:multi, or we can ask one or more labelers to provide each label separately, as in the right panel (and also in Figure 1).
The choice between these two approaches echos the choice in surveys between select-all-that-apply questions and a series of yes/no questions (Figure LABEL:fig:selectall). As predicted by the survey response model, the yes/no format collects better data because it encourages respondents to process each option separately (Smyth, 2006; Pew Research Center, 2019). The select-all-that-apply approach is vulnerable to satisficing: respondents pick the first one or two reasonable options and fail to think deeply or even look at later options (Galesic et al., 2008). However, the yes/no format can encourage acquiescence (Smyth, 2006).
This finding also holds for labeling: when collecting two labels of the same observation, splitting the collection across two screens (the left panel of Figure LABEL:fig:multi) results in better performing models (Kern et al., 2023). This result is a clear example of how findings in the survey literature translate to the labeling task and improve the quality of training data.
@@
The inclusion of ‘don’t know’ or ‘no opinion’ responses in surveys has been debated for years. Some researchers believe these options offer respondents an easy way to satisfice: rather than thinking about the issue and forming an opinion, respondents can simply choose the ‘don’t know’ option. Others believe that having no opinion on a given topic is a valid response and that forcing respondents to provide an opinion when they don’t have one reduces data quality (Schuman & Presser, 1996).
Many labeling tasks do not include a ‘skip’ or ‘don’t know’ option. Labelers must provide a label even when they are not certain. When a recent experiment provided a ‘don’t know’ option to half of the labelers, fewer than three percent chose it, and the overall distribution of the labels was not impacted (Beck et al., 2022). Another recent study in NLP collected ‘uncertain’ flags from labelers for a relation extraction task across different text genres Bassignana & Plank (2022). Some genres were more challenging to label, and the classifier also struggled with these items. These preliminary results suggest that giving labelers the option to indicate uncertainty or lack of knowledge provides helpful additional information and does not encourage satisficing in labeling.
@@
Pre-labeling involves displaying a suggested label, bounding box, or similar and asking the labeler if it is correct. If the labeler indicates the label is not correct, they are asked to provide the correct label. Pre-labeling is more efficient than labeling without suggested labels (Lingren et al., 2013; Skeppstedt et al., 2016; South et al., 2014). However, labelers may become too trusting of model suggestions and fail to correct errors (Dietvorst et al., 2015; Logg, 2017; Berzak et al., 2016), a phenomenon called anchoring bias or automation bias (Mosier & Skitka, 1999).
@@
Previous literature on labeling has explored anchoring bias (Lingren et al., 2013; Skeppstedt et al., 2016; South et al., 2014) but has not leveraged social science to find the factors that make make the effect weaker or stronger. The social science literature suggests several hypotheses about the mechanisms behind anchoring bias, such as incentives (Cialdini, 2009), belief in authority (Asch, 2016; Cialdini, 2009), or reliance on heuristics (Cialdini, 2009; Norman, 2007; Kahneman, 2011). Testing these theories experimentally would help design tasks that capture the efficiency of pre-labeling with lower risk of anchoring bias.
@@
Biases similar to anchoring from pre-labeling can evolve through examples. Survey questions often give examples of the things respondents should consider when they formulate their responses. Examples improve response accuracy when they remind respondents to include items they might otherwise leave out, because they have forgotten or were unsure whether to include them. However, when the examples include only frequent items, respondents tend to leave out infrequent items more often than they did before the examples were provided (Tourangeau et al., 2014).
Labeling instructions or guidelines also often include examples. As in surveys, labelers at times rely too heavily on these examples as they label. This instruction bias can lead to overestimation of model performance (Parmar et al., 2023).

### v2

# Position: Insights from Survey Methodology can Improve Training Data

@@ Abstract
Whether future AI models are fair, trustworthy, and aligned with the public’s interests rests in part on our ability to collect accurate data about what we want the models to do. However, collecting high-quality data is difficult, and few AI/ML researchers are trained in data collection methods. Recent research in data-centric AI has show that higher quality training data leads to better performing models, making this the right moment to introduce AI/ML researchers to the field of survey methodology, the science of data collection. We summarize insights from the survey methodology literature and discuss how they can improve the quality of training and feedback data. We also suggest collaborative research ideas into how biases in data collection can be mitigated, making models more accurate and human-centric.
Training Data, Data Quality, Surveys, Noisy Labels, Annotation
@@ 1 Introduction
Social scientists have long relied on survey data collected from human subjects to quantify the population, understand public opinion, and test hypotheses about human behavior. The methods used to collect survey data have been extensively studied and refined by researchers in the field of survey methodology, which draws on social and cognitive psychology to develop theories about how humans understand, process, and respond to questions in surveys (Groves et al., 2009).
Data labeled by humans is also central to all stages of the AI pipeline (Plank, 2022; Mazumder et al., 2023), from initial model training, to fine-tuning, reinforcement learning, and model assessment. Insights from social science can contribute to the development of more trustworthy and human-centric models: “if we want to train AI to do what humans want, we need to study humans” (Irving & Askell, 2019). However, collecting high-quality data is difficult, as decades of research in survey methodology and recent high-profile failures in opinion polling (Sturgis et al., 2016; Kennedy et al., 2017; Clinton et al., 2021) demonstrate.
Given the importance of human-labeled data to AI model development, we are surprised that little research in the AI literature has used social science, and survey methodology in particular, to understand the actions and motivations of the humans behind the data generating process. We worry that many researchers collecting data to train, fine-tune, or reinforce AI and ML models are not trained in data collection. A recent paper lamented that, among AI researchers, “everyone wants to do the model work, not the data work” (Sambasivan et al., 2021).
This position paper argues that lessons from survey methodology can improve the quality and efficiency of training data and thus improve models trained on those data. We introduce AI researchers to the community of scientists who want to do the data work and their insights into how to collect high-quality data. We first make the case that label collection is similar to survey data collection (Section 2). Next, we draw on social science theories to develop hypotheses about the facets of the data collection task that may impact the quality of the labels collected (Section 3). Then, we discuss who works as labelers and how the characteristics and uniqueness of the labelers can impact the labels collected and the models trained on those data (Section 4). In Section 5, we join the call for greater transparency in label collection methods, offering lessons from surveys and statistical methods. Throughout the paper, we use the terms labels (and labelers) to refer generally to ML annotations, such as image object labels and bounding boxes, natural language understanding labels, model evaluation data, human feedback for reinforcement learning, and other types of training data (and data generators).
@@ 2 How Labeling is Like a Survey
Surveys can be self-administered, like a web or paper-and-pencil survey, or interviewer administered, like a telephone or face-to-face survey. A web survey presents a series of questions with, most often, closed answer choices. Superficially, the process of labeling observations for ML models often looks like a web survey: labelers see one or more prompts and associated answer choices (Figure 1). As in surveys, the task can be factual (is there a bicycle in this photo?) or opinion based (which of these responses to the prompt is most helpful?).
@@
Of course, label collection differs from survey data collection in important ways. Surveys ask many different questions and end when the questions run out; labeling tasks usually consist of similar repeated observations, and labelers often continue until they choose to stop. Surveys usually ask people about themselves: their opinions, behaviors, characteristics. Labeling tasks more often involve passing judgment on an object outside of oneself (for example, images, product reviews, or a news article). However, the human feedback data used in reinforcement learning often aims to capture personal opinions: “which of the following responses is most relevant to the prompt?”
@@
The goal of surveys and labeling tasks is also different. Surveys ask questions of a sample of selected persons to make inference about the population. For example, say we wished to know what proportion of the U.S. adult population does not have health insurance. We select a sample of U.S. adults and ask them whether they have health insurance. For each respondent, we have a yes or no response to the question. [^1] We are ignoring “I don’t know” and other types of responses. The proportion of respondents without insurance in the sample is an estimate of the proportion of the population without insurance. A survey often asks many questions on related topics. Our example survey might also ask “when was the last time you visited a health care professional?” and “have you received a flu vaccine in the last 12 months?” Results are often reported at the question level although relationships between questions are also of interest: for example, are people with health insurance more likely to get a flu vaccine? The quality of the analysis rests on collecting accurate responses from a representative sample.
@@
The goal when collecting labeled data is not to estimate the proportion of the population that finds a given post offensive or detects a vehicle in a given image. Instead, the goal is to learn patterns from the labeled data to predict labels for unseen observations. Thus, while it is not as important that labels reflect the population’s views at the observation level, across observations, the views of the population should be represented. Collecting accurate labels from a diverse set of labelers is important to the performance and generalizability of the final model.
Despite their differences, both surveys and label collection have these two needs in common. The data points provided in response to a prompt should capture the data provider’s judgment: accurate responses / accurate labels. Those who provide the data points should represent the judgments of the relevant population: representative sample / diverse set of labelers. Sections 3 and 4 discuss these needs in turn.
3 Need for Accuracy
Survey methodologists have developed theories of how respondents understand, process, and respond to survey questions. We summarize these theories and use them to derive hypotheses about the aspects of the label collection task that may impact label quality. We end the section with thoughts on mitigation measures and future research.
3.1 Response Process
Ideally, survey respondents understand questions thoroughly and respond thoughtfully. However, they make take shortcuts that can threaten data quality.
@@ 3.1 Response Process
Responding to a survey question can involve several cognitive steps (Tourangeau et al., 2000; Tourangeau, 2018):
@@
Ideally, a respondent proceeds through each step in order. However, they can choose to backtrack. For example, considering the response options in the Mapping step may change the interpretation of the question (Comprehension) or trigger additional relevant information to come to mind (Retrieval).
The above model exposes why respondents sometimes give incorrect answers. At the Comprehension step, they may fail to understand the question or some of the words it uses. They may have a different understanding of some of the words than those who wrote the question. At the second step, respondents may fail to retrieve all relevant information. Some information may have been forgotten. At the third step, respondents may fail to put in the mental effort to bring together their understanding of the question with the retrieved information. At the fourth step, respondents may not find an answer choice that reflects their answer or they may edit the true answer to avoid revealing sensitive information.
@@
The full survey response process outlined above is cognitively demanding. Some respondents resort to taking shortcuts, an approach called satisficing (Krosnick, 1991; Krosnick et al., 1996). For example, they may retrieve only the most recent relevant information from memory (recency bias) or choose the first reasonably correct answer choice. Satisficing relates to the cognitive miser theory in psychology, which holds that people seek to minimize cognitive effort (Fiske & Taylor, 1991; Kahneman, 2011).
As predicted by satisficing theory, eye-tracking studies show that respondents do not read all options in select-all-that-apply questions (Galesic et al., 2008), and shortcuts are more common as survey length increases (Galesic & Bosnjak, 2009). Respondents tend not to read provided instructions (Brosnan et al., 2019) or click on provided definitions (Peytchev et al., 2010), especially when they believe they understand the concept that is asked about (Tourangeau et al., 2006).
The survey literature discusses several types of more extreme undesirable response behavior. Acquiescence is the tendency to say “yes” to yes/no questions, regardless of content (Knowles & Condon, 1999). Straightlining is the practice of choosing the same response option in the same position (for example, the first response option) to all questions. This behavior is most common in batteries or grids of questions with the same response options (Kim et al., 2019). Some respondents even deliberately give incorrect answers to later questions to reduce the length or burden of a survey: when a “yes” response triggers follow up questions, respondents may learn to report “no.” This phenomenon is called motivated misreporting (Kreuter et al., 2011; Tourangeau et al., 2012; Eckman et al., 2014).
@@
Perceptions and judgments are shaped by the broader context and preceding experiences, a phenomenon called context effects (Tversky & Kahneman, 1974; Strack, 1992). For example, a very tall person can make others seem shorter: a contrast effect. An unethical politician can make other politicians seem less ethical: an assimilation effect (Bless & Schwarz, 2010).
Opinion questions are especially vulnerable to context effects, because respondents do not always have well-formed, fixed opinions that they retrieve from memory. Instead, they form opinions when asked for them, and this process can be shaped by context clues in the question, the response options, the look and feel of the instrument, or the previous questions (McFarland, 1981; Zaller & Feldman, 1992; Schwarz, 2007).
Order effects are the most common example of context effects in surveys. Questions that come earlier in a survey can change how respondents interpret later questions. Researchers have found order effects in reports of crime and bullying victimizations, of disabilities, and of race (Cowan et al., 1978; Gibson et al., 1978; Bates et al., 1995; Todorov, 2000; Huang & Cornell, 2015).
@@ 3.2 Hypotheses about Label Quality
These theories about how respondents answer questions lead us to several hypotheses about the properties of the labeling task that may impact training data quality. The ML literature has investigated some of these hypotheses, but fundamental research gaps exist.
Wording and Reading Level
Respondents cannot provide high-quality answers to questions unless they clearly understand what the question asks and what they should include and exclude in their answer. Questions should be at an eighth grade reading level or lower (Dillman et al., 2014) and terms should be as unambiguous as possible. Definitions, if needed, should be provided in the question itself, because respondents often do not use rollovers or links for additional information when answering questions (Peytchev et al., 2010).
We suspect that applying the same guidelines to labeler prompts and instructions would improve the quality of the labels collected. (Of course, nothing can be done about ambiguous terms or high reading levels in the observations.) We are not aware of any research into these issues. We note that Figure 1 contains rollovers or links for additional information on three of the questions on the right side.
@@
Often ML researchers want to collect multiple labels about an observation from one labeler: for example, whether an image contains a cat, a dog, a person, or a vehicle. We can ask labelers to provide all labels at once, as in Figure LABEL:fig:1screen, or we can ask one or more labelers to provide each label separately, as in Figure LABEL:fig:sep_screens.
The choice between these two approaches echos the choice in surveys between select-all-that-apply questions (Figure LABEL:fig:select_all) and a series of yes/no questions (Figure LABEL:fig:yesno). As predicted by the survey response model, the yes/no format collects better data because it encourages respondents to process each option separately (Smyth, 2006; Pew Research Center, 2019). The select-all-that-apply approach is vulnerable to satisficing: respondents pick the first one or two reasonable options and fail to think deeply or even look at later options (Galesic et al., 2008). However, the yes/no format can encourage acquiescence (Smyth, 2006).
This finding also holds for labeling. In an experiment that involved labeling tweets as containing hate speech or offensive language, Kern et al. (2023) randomly assigned labelers to different versions of the labeling instrument. Condition A was similar to Figure LABEL:fig:1screen and Condition B was similar to Figure LABEL:fig:sep_screens. Splitting the collection across two screens (Condition B) led to higher rates of hate speech and offensive language annotation. Models trained on Condition B data also performed better than those trained on Condition A data across several metrics (Kern et al., 2023). This result is a clear example of how findings in the survey literature translate to the labeling task and improve the quality of training data.
Order Effects
Theories about context effects suggest that the order in which instances or observations are presented influences the labels assigned. If a contrast effect is present, a very hateful social media post would make later posts seem less hateful than they otherwise would (for preliminary evidence of this phenomenon, see Beck et al., 2022). An order effect could also arise if labelers change their behavior over time. As they gain experience, they might become more accurate and conscientious, as suggested by Lee et al. (2022). Alternatively, they might become bored or fatigued and engage in more satisficing, acquiescence, or “don’t know” nonresponding over time, as suggested by the survey literature (Kraut et al., 1975; Galesic & Bosnjak, 2009).
Mathur et al. (2017) detected order effects in two benchmark NLP data sets. When coding tweets as hate speech or offensive language, Beck et al. (2024) found a decreasing time trend: tweets that appeared later were less likely to be flagged. However, their experimental set up did not allow them to test hypotheses about the mechanisms causing the time trend.
Many research questions thus remain open, such as when contrast and assimilation effects appear and which tasks and labelers are most impacted by order effects. Order effects may also have implications for active learning (AL) and similar labeling approaches. In AL, an algorithm determines which observations to label next to maximize the marginal information gain for the model (Monarch, 2021). However, active learning considers only the model’s needs, not the labelers’. If context effects are large, the algorithm should also account for contrast and assimilation effects when deciding which observations to label. We are not aware of any research that has jointly accounted for the needs of the model for diverse training data and the impact of observation order on annotators. An approach similar to active learning exists in surveys (Zhang et al., 2020), but is not widely adopted due in part to concerns about order effects.
@@
The inclusion of “don’t know” or “no opinion” responses in surveys has been debated for years. Some researchers believe these options offer respondents an easy way to satisfice: rather than thinking about the issue and forming an opinion, respondents can simply choose the “don’t know” option. Others believe that having no opinion on a given topic is a valid response and that forcing respondents to provide an opinion when they don’t have one reduces data quality (Schuman & Presser, 1996).
Many labeling tasks do not include a “skip” or “don’t know” option: labelers must provide a label even when they are not certain. (The instrument in Figure 1 is an exception.) When a recent experiment provided a “don’t know” option to half of the labelers, fewer than three percent chose it, and the overall distribution of the labels was not impacted (Beck et al., 2022). Another recent study in NLP collected “uncertain” flags from labelers for a relation extraction task across several text genres (Bassignana & Plank, 2022). Labelers were more likely to choose “uncertain” when coding text in some genres, and the model struggled with prediction in those genres as well. These preliminary results suggest that giving labelers the option to indicate uncertainty or lack of knowledge can provide helpful information and does not encourage satisficing in labeling.
@@
Pre-labeling involves displaying a suggested label, bounding box, or similar and asking the labeler if it is correct. If the labeler indicates the label is not correct, they are asked to provide the correct label. Pre-labeling is more efficient than labeling without suggested labels (Lingren et al., 2013; Skeppstedt et al., 2016; South et al., 2014). However, labelers may become too trusting of the suggestions and fail to correct errors (Dietvorst et al., 2015; Logg, 2017; Berzak et al., 2016), a phenomenon called anchoring bias or automation bias (Mosier & Skitka, 1999).
@@
Previous literature on labeling has explored anchoring bias (Lingren et al., 2013; Skeppstedt et al., 2016; South et al., 2014) but has not leveraged social science to find the factors that make the effect weaker or stronger. The social science literature suggests several hypotheses about the mechanisms behind anchoring bias, such as incentives (Cialdini, 2009), belief in authority (Asch, 2016; Cialdini, 2009), or reliance on heuristics (Cialdini, 2009; Norman, 2007; Kahneman, 2011). Testing these theories experimentally would help data collectors design tasks that capture the efficiency of pre-labeling with lower risk of anchoring bias.
@@
Examples can introduce a similar bias. Survey questions often give examples of the things respondents should consider when they formulate their responses. Examples improve response accuracy when they remind respondents to include items they might otherwise leave out, because they have forgotten or were unsure whether to include them. However, when the examples include only common items, respondents tend to leave out less common items (Tourangeau et al., 2014).
Examples are also often included in labeling instructions or annotation guidelines. As in surveys, labelers at times rely too heavily on these examples as they label. This instruction bias can lead to overestimation of model performance (Parmar et al., 2023). Again, a better understanding of the mechanisms behind this behavior, guided by social science theories, could inform efforts to reduce it.